OpenStudy (aristasiebrits8):

Hi! i need a bit of help on this question. Medal for help! Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). Options: y = x + 7 y = -3x + 3 y = x + 3 y = -3x + 7 Thanks so much!

OpenStudy (geliraymundo):

From the equation we can see that the slope is -3 which will also be the slope for the other equation and we'll use y-y1 = m(x-x1) so it'll be y- 6 = -3(x-(-1)). Simplifying you'll have y-6 = -3x-3. Simplify some more we'll have y = -3x+3

OpenStudy (aristasiebrits8):

Thanks so much! Can you maybe help me with another one?

OpenStudy (geliraymundo):

Well, I hope I can answer it.

OpenStudy (aristasiebrits8):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = - x - 6 y = - x - 14 y = 3x - 6 y = 3x - 14

OpenStudy (geliraymundo):

Start by making the equation y=x+b. So, the given equation will be y=3x+3. From that we know that the slope is 3. Same with earlier with use the point-slope form, so it'll become y-(-2)=3(x-4). Simplifying will make it y+2=3x-12. Simplifying it further will make it y=3x-14

OpenStudy (aristasiebrits8):

Okay! Thanks for the explanation as well! I know it is a lot to ask, but I only have two more questions left. Would you mind helping me on those as well?

OpenStudy (geliraymundo):

Basically, for parallel copy the coefficient of x and that is your slope and then substitute the points given. For perpendicular you need to get the negative reciprocal. What if you try answering and maybe I can check so that you'll have a practice with those kinds of questions. Deal?

OpenStudy (aristasiebrits8):

Sure!

OpenStudy (aristasiebrits8):

So first question: Select the equation of a line that is perpendicular to the line on the graph and passes through the point (3, 2). (1 point) y =1/3 x + 3 y = -1/3x+3 y = 3 x + 2 y = -1/3 x + 2 I think it is the third one, because it includes both of the coordinates in the point.

OpenStudy (phi):

***I think it is the third one, because it includes both of the coordinates in the point.*** that shows you are thinking.. but you need to use different info. You need to know the slope of the line. the slope is the number "in front" of the x (i.e. that multiplies the x) in this problem, you also need to know the slope of a perpendicular line

OpenStudy (aristasiebrits8):

Ok. How do I know that?

OpenStudy (phi):

you need the slope of ** perpendicular to the line on the graph ** first step: find the slope of the line in the graph. find a "nice point" where the line goes through two lines that cross. then count up to the next place the line goes through two crossing lines, then how far over. divide change in y divided by change in x to get the slope

OpenStudy (phi):

for example, put your finger on the point at (-1,-1) can you find another point where the line goes through two crossing lines. ?

OpenStudy (aristasiebrits8):

it crosses through -2 if you go left down

OpenStudy (phi):

you want a spot where it crosses both an x and y line (so that we can count)

OpenStudy (geliraymundo):

It's actually the first. You forgot to look for the slope... and get the negative reciprocal. Looking at the graph you could see that there is a point with coordinates (1,5) and another with (0,2). So, you can compute the slope by getting the difference of y over difference of x. So, here we can use 5-2/1-0 which is equal to 3. A technique is label the points as P1 and P2 and if you choose the y to be P2-P1 the x should be in that order too. then to get the negative reciprocal multiply it with negative 1 and interchange the denominator and numerator. In this case the denominator is 1 making it -1/3. Then just like the one with the parallel lines, substitute the given points it has to cross.

OpenStudy (phi):

in other words, the "corner" of a "box"

OpenStudy (geliraymundo):

I don't know if my explanation of slope made sense but think of slope as rise over run and if you're a person who's into formulas it's (y2-y1)/(x2-x1). Tell me if I should explain it more.

OpenStudy (aristasiebrits8):

Okay! Now I understand.

OpenStudy (aristasiebrits8):

So for question 2... Select the equation of the line that passes through the point (-2, 3) and is perpendicular to the line on the graph. y = 0 y = -2 y = 3 y = 3x

OpenStudy (aristasiebrits8):

And I think it is y=3

OpenStudy (geliraymundo):

The line in the graph is vertical giving it a undefined slope. For a line to be perpendicular to it, it has to be horizontal. Horizontal lines have a slope of 0. So, the equation is y-3=0. And yes you're correct it's y=3. Good job!